The mass of a sample bottle is 27.23g. When filled with water, the bottle weighs 46.85g. The density of water is known to be 0.9976 g/mL. Calculate the capacity (volume) of the bottle.
1
Expert's answer
2018-02-07T04:53:19-0500
(bоt.) = 27.23 g
m( bоt. + H2O) =46.85g
ρ(H2O) = 0.9976 g/ ml
V(bоt.) - ?
V(bоt.) = V(H2O)
V(H2O) = m(H2O)/ ρ(H2O) =( m( bоt. + H2O) - m(bоt.))/ ρ(H2O) = (46.85g - 27.23 g)/0.9976 g/ml = = 19.67 ml